Can you help me with this problem please! : The length of a rectangle is 4 cm less that three times its width. If the length is increased by 8 cm and the width is decreased by 2 cm, the perimeter will be 100 cm. Find the dimensions of the original rectangle.
Answer by Nate(3500)

Original:
width = w
length = 3w - 4
Tampered:
width = w - 2
length = 3w + 4
2l + 2w = perimeter
2(3w + 4) + 2(w - 2) = 100
6w + 8 + 2w - 4 = 100
8w = 96
w = 12
width = 12 - 2 = 10cm
length = 3(12) + 4 = 36 + 4 = 40cm
Original:
width = w = 12cm
length = 3(12) - 4 = 36 - 4 = 32cm